Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-2015-03-27

Observer : Nic Walton, Heather Campbell
Support team : Southampton (Mat Smith and Régis Cartier)
TAT :

NOTE: Data transfer problems due to issues with the connection between the observing and data reduction machines (roll over problem from earlier powercuts). Support staff not able to fix. Hence was not possible to analyse the data during the night.  Thus Atel will need to be done next night for the classification targets below. 

Start of night: Clouds, rain, thunder and lightening!

End of night: 

Seeing: Measured ??? at start of night, ??? for most of night.

Afternoon calibrations:  EFOSC arcs and specflats

22.55    Sunset

Closed at start of night

02.10 Opened telescope!

checking telescope working (as not been opened for 4 nights) and image quality first

02.28 AS15fr spectra (classification)


02.35 PSNJ07361576 or SN2015F GR16+GR11


03.10 OGLE173 spectra (classification)


03.25 PS15si spectra (classification)


03.40 PSNJ10544216-2104138


seeing ~1.5


06.00 PS15ae follow up G11+G16 with slit width 1.5'


07.15 PSN15br G13 only 1 as too close to the limit


08.00 LTT7379 standard


08.25 LSQ15rw G13


09.00 SNHunt G13+BVRi imaging


09.20 PS15rs spectra (classification)


09.35 PS15vs spectra (classification)


09.45 EG274 standard


10.00PS15oz - trying but might be too bright (classification)


10.25 too bright for skyflats







Comments